OpenStudy (anonymous):

I WILL FAN AND MEDAL! Find the product of (x + 5)(x − 5). x2 − 10x + 25 x2 + 10x + 25 x2 − 25 x2 + 25

OpenStudy (anonymous):

3rd option

OpenStudy (anonymous):

Doing it in my head, I believe it ends up being difference of squares, which is the third option

OpenStudy (anonymous):

Its a difference of squares

hero (hero):

So much for Hints

OpenStudy (anonymous):

Third option

OpenStudy (anonymous):

guess what? \(\huge \color{red}C, \color{blue}C, \color{green}C\)

OpenStudy (anonymous):

again it is \((a+b)(a-b)=a^2-b^2\)

OpenStudy (anonymous):

LMAOOOOOOOOOOOOO XD U have such a great humor @satellite73

OpenStudy (anonymous):

\(\text{by the }\color{red}C,\text{by the } \color{blue}C,\text{by the beautiful } \color{green}C\)

OpenStudy (anonymous):

Lmaooooo U got me in tears

hero (hero):

\[\begin{array}l\color{red}{\normalsize\text{I}}\color{orange}{\normalsize\text{ }}\color{#9c9a2e}{\normalsize\text{k}}\color{green}{\normalsize\text{i}}\color{blue}{\normalsize\text{n}}\color{purple}{\normalsize\text{d}}\color{purple}{\normalsize\text{a}}\color{red}{\normalsize\text{ }}\color{orange}{\normalsize\text{s}}\color{#9c9a2e}{\normalsize\text{t}}\color{green}{\normalsize\text{i}}\color{blue}{\normalsize\text{l}}\color{purple}{\normalsize\text{l}}\color{purple}{\normalsize\text{ }}\color{red}{\normalsize\text{l}}\color{orange}{\normalsize\text{i}}\color{#9c9a2e}{\normalsize\text{k}}\color{green}{\normalsize\text{e}}\color{blue}{\normalsize\text{ }}\color{purple}{\normalsize\text{t}}\color{purple}{\normalsize\text{h}}\color{red}{\normalsize\text{i}}\color{orange}{\normalsize\text{s}}\color{#9c9a2e}{\normalsize\text{}}\end{array}\]